OpenStudy (anonymous):

Consider two point like objects with excess charge on them. Suppose they lie along the x-axis. A net charge of 2.0x10^-9 C is locates x=3.0 cm and a net charge of -3.0x10^-9 C point charge is located at x=5.0 cm. What is the magnitude of the force on the negatively charged object due to the presence of the positively charged objects? What is its direction? Express the force as a vector quantity using unit vector notation.

OpenStudy (lastdaywork):

Would you like to make an attempt ??

OpenStudy (anonymous):

k= 9.0x10^9 Nm^2/C^2 q1= 2.0x10^-9 C ? q2= -3.0x10^-9 C ? so i need to find the distance between 3.0 cm and 5.0 cm ?????

OpenStudy (lastdaywork):

Yes..keep going :)

OpenStudy (anonymous):

F= (k.q.q)/(r^2) = ((9.0x10^9)(2.0x10^-9)(-3.0x10^-9))/(2^2)=-1.35 × 10^-8

OpenStudy (lastdaywork):

Personally; I'll advice you to only find magnitude by calculations and use physics to determine the sign. For example; in this case we have two unlike charges, hence attraction. So the force vector would be directed towards the other charge.

OpenStudy (anonymous):

ok thanks.d

OpenStudy (lastdaywork):

:)